# What You Say = What You Want? Teaching Humans to Articulate Requirements for LLMs

### v1

@@ Abstract.
Prompting ChatGPT to achieve complex goals (e.g., creating a customer support chatbot) often demands meticulous prompt engineering, including aspects like fluent writing and chain-of-thought techniques. While emerging prompt optimizers can automatically refine many of these aspects, we argue that clearly conveying customized requirements (e.g., how to handle diverse inputs) remains a human-centric challenge. In this work, we introduce Requirement-Oriented Prompt Engineering (ROPE), a paradigm that focuses human attention on generating clear, complete requirements during prompting. We implement ROPE through an assessment and training suite that provides deliberate practice with LLM-generated feedback. In a study with 30 novices, we show that requirement-focused training doubles novices’ prompting performance, significantly outperforming conventional prompt engineering training and prompt optimization. We also demonstrate that high-quality LLM outputs are directly tied to the quality of input requirements. Our work paves the way for more effective task delegation in human-LLM collaborative prompting.
LLM, prompt engineering, requirement engineering, end-user programming
@@
Figure 1. We propose Requirement-Oriented-Prompt Engineering (ROPE) to help novices write effective prompt programs by focusing on articulating good requirements. (A) Our training provides deliberate practice in refining requirements, with automated feedback across various modalities. (B) We provide assessments that link requirement quality to LLM output and evaluate on these assessments, in a counter-balanced pre-post experimental design. (C) Through pre-post assessments where novices write prompts to replicate existing programs (e.g. replicating Trip Advisor in D through a prompt in E), we observe that ROPE significantly improves novices’ prompts and requirements compared to traditional prompt engineering training.
@@ 1. Introduction
General-purpose AIs like large-language models (LLMs) have evolved from simple next-word predictors (Brown, 2020) to powerful assistants capable of fulfilling complex user needs (Qin et al., 2024; He et al., 2024; Ouyang et al., 2022). This improvement in LLM instruction-following has encouraged users to delegate increasingly intricate tasks to these models. In the early days of prompt engineering, users primarily focused on refining the wording of simple instructions to improve LLM output quality (Jiang et al., 2020; Shin et al., 2020). Today, prompts resemble detailed “essays” that define LLM roles, human preferences, and other task-specific requirements. Rather than one-off small requests, these prompts start to power the generation of programs. On one hand, designers and developers now write functional descriptions for LLM agents (e.g., Devin, SWE-Agent) to translate into executable software code (Wodecki, 2024; Shani, 2023; Jimenez et al., 2024; Yang et al., 2024a). On the other hand, everyday users can write complex prompts to tailor general-purpose LLMs into special purpose LLM Applications. For instance, an LLM app (or a GPTs) like Trip Advisor  [^1] https://chatgpt.com/g/g-K38J9feSL-trip-advisor (Figure 1D) can be built solely using prompts similar to Figure 1E. These LLM applications function similarly to traditional software and are accessible through LLM App Stores like GPT Store (Unk, [n. d.]) and FlowGPT (FlowGPT, [n. d.]), where they can be reused with a single click. In just a few months, over 30,000 “LLM app developers” have created millions of GPT apps, with the most popular ones being used more than 5 million times (Su et al., 2024; Zhang et al., 2024). This trend signals a future of end-user programming through natural language, where anyone can create reusable programs just by writing prompts.
@@
The future of writing prompt programs looks promising, but status quo prompt engineering remains a challenge. Various studies show that optimizing LLM outputs requires well-rounded prompts, involving fluent writing, clear instructions, personas, formats, and effective adaption of prompting techniques like Chain-of-Thought (Bsharat et al., 2023; Meskó, 2023; Marvin et al., 2024; Schmidt et al., 2024; Desmond and Brachman, 2024; freeCodeCamp.org, 2023). These complexities often confuse novices, causing them to make ad hoc revisions to prompts without a clear understanding of what needs improvement (Zamfirescu-Pereira et al., 2023). Recognizing the challenge, the NLP community is developing prompt optimizers to automate some prompt refinement steps, e.g., incorporating role-plays, applying effective techniques, structuring prompts, and enhancing text fluency (e.g., Figure 2, discussions in Section 2.2).
@@
The rise of optimizers makes us think: What aspects of prompt engineering should humans still master? We argue that the remaining part of manual prompt creation is akin to the core step before programming: requirement articulation, where natural language requirements are used to define and document how a system should behave, including expected inputs and outputs (Lamsweerde, 2009) (e.g., ‘‘Ask follow-ups’’ for Trip Advisor app in Figure 1). Requirements are the backbone of program functionalities. The prompt in Figure 2 drives a complex LLM adaptation because the user manually provided specific requirements. Customized requirements are hard to predict, and thus difficult to fully automate and must be written by humans. We argue that in a future where prompt programs are common, individuals should develop proficiency in requirement articulation. In other words, we need to shift our focus towards Requirement-Oriented Prompt Engineering, a novel paradigm that we call “ROPE.
@@
Articulating clear and complete requirements is a known challenge in software engineering (Liu et al., 2023b; Nam et al., 2023; Nguyen et al., 2024), and poor requirements often cause software failures (Millett et al., 2007). This difficulty is amplified by the unpredictable nature of LLMs, which complicates controlling outputs (Denny et al., 2024; Chopra et al., 2023). Even experts need multiple iterations to refine requirements (Shankar et al., 2024a), and novices frequently make requirement mistakes (Babe et al., 2023). Recognizing both the importance and difficulty of this task, we pose the research question: Can we help end-users better instruct LLMs to achieve their goals through ROPE training?
@@
Figure 2. Automatically refine a user’s prompt for Trip Advisor using the optimizer Prompt Maker. Customized requirements still need to come from the human prompt, while the optimizer improves on fluency, role plays, structures, etc.
@@
To answer this question, we develop the ROPE paradigm (Section 3), along with the requirement-focused training and assessments to (1) help users practice writing complete and correct requirements, and (2) deepen our understanding of how requirements affect the quality of LLM outputs. Specifically, we formalize tasks (Section 4.1) with complex requirements (e.g., game development), and design corresponding assessments (Section 4.2) that can measure the requirement quality, the LLM output quality, and their correlations. We also design and implement an interactive training mechanism (Section 4.3) that disentangles requirement articulation from other adjacent prompting aspects. Our system supports deliberate practice on requirement refinement by generating immediate feedback grounded on oracle requirements.
@@
We conduct a randomized controlled user study (Section 5) with 30 prompting novices, comparing the requirements users write before and after training, and against a control group that received standard prompt engineering training. We confirm that our requirement-focused training is effective (Section 6): novices receiving this training significantly improved their requirement performances by two-fold in the pre-to-post test, achieved significantly more gains than the control group, and learned to iterate on prompts in a more structured way towards requirements. Moreover, we confirm that requirements are indeed important for the LLM generation: There is a strong positive correlation between the requirement qualities and the LLM output qualities. Importantly, the training gains are complementary to potential improvements from optimizers: Using a popular prompt optimizer, we observe that optimizers can indeed improve requirement quality, but they cannot close the gap between the baseline and the training groups. We conclude with an in-depth discussion (Section 7) on future directions for human training and optimizer design to enhance human-LLM complementarity in ROPE. In summary, we contribute:
@@
As natural language prompt programs become more widespread, everyone needs the skill to write good requirements. We believe that with our ROPE paradigm, LLMs can empower anyone to program reusable AI assistants.
@@ 2. Related Works
We review three key areas relevant to our work: prompt engineering, prompt optimization in natural language processing, and requirement engineering in end-user software engineering. We highlight a significant gap in the literature: the lack of requirement-oriented prompt engineering training for end users, which we propose to address in this paper.
@@ 2.1. Current Prompt Engineering Practices and Challenges
Prompt engineering (PE) has been essential for crafting effective input instructions to guide LLMs toward generating desired outputs (Liu et al., 2021; Bach et al., 2022; Zamfirescu-Pereira et al., 2023). However, the non-deterministic nature of LLMs makes it challenging for humans to predict LLMs’ behavior across prompts (Ouyang et al., 2023; Denny et al., 2024). This leads to wasted time on unproductive strategies, such as trivial wording changes (Feldman and Anderson, 2024; Denny et al., 2024; Zamfirescu-Pereira et al., 2023). While some challenges can be mitigated through automation (see Section 2.2), e.g., automating word choices, human requirements remains crucial for customized or specialized tasks (Liu et al., 2021). We define these task as LLM-hard, as a simple prompt cannot produce a satisfactory response. Developing reusable prompts for customized chatbots or GPTs (Bach et al., 2022; Zamfirescu-Pereira et al., 2023) is a common LLM-hard task. While a prompt is an input to LLMs, a reusable prompt is a set of instructions for recurring goals, essentially functioning as a prompt program. As more users create reusable prompts, the ability to write good prompt programs with clear requirements becomes increasingly important.
@@
However, prompt creation can be complex and inaccessible to non-experts who lack technical knowledge of LLMs, highlighting the need for more end-user prompt engineering (EUPE) scaffolding and training (Zamfirescu-Pereira et al., 2023). Various strategies have been developed to support PE, such as prompt technique catalog (Bsharat et al., 2023), chat-based tools like GPT-builder, [^2] https://chatgpt.com/create and toolkit for orchestrating complex tasks (Chase, 2022; Wu et al., 2024; Khattab et al., 2023). Each tool has trade-offs, balancing flexibility, precision, and technicality. For example, chat-based tools may ask narrow clarifications without fully understanding the context, leaving users to self-refine requirements when LLM responses are open-ended (Lahiri et al., 2022; Qian et al., 2024).
@@
Existing PE training emphasizes the importance of explicitly stating requirements within prompts (Meskó, 2023; Schmidt et al., 2024; Desmond and Brachman, 2024), yet there is still a lack of focused training on requirement generation skills for everyday users. Without training, non-experts often make mistakes such as missing requirements and writing conflicting requirements in prompts (Feldman and Anderson, 2024; Nam et al., 2023; Denny et al., 2024). Even experts need many iterations to improve their requirements for complicated LLM tasks (Shankar et al., 2024a). However, training novices on prompt engineering is difficult; for instance, novice programmers may not improve at prompting within a 75 minute study when provided with test cases and generated code as feedback (Nguyen et al., 2024). We aim to address this gap by proposing a ROPE training for end-users to improve their requirement engineering skills in the context of prompt creation.
@@ 2.2. Instruction Following for Foundation Models
Optimizing LLM performance on customized tasks typically follows two approaches: improving the models directly, or refining prompts with models fixed. The former has enabled the versatility of models like GPT-4 (OpenAI, 2023), Gemini (DeepMind, 2023), and LLaMA 3 (AI, 2023), which are fine-tuned through instruction tuning (Ouyang et al., 2022) and further aligned with human preferences using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). For the latter, the NLP community has been exploring LLMs’ capabilities to follow requirements (or “constraints”). Various datasets have been proposed to assess whether LLMs can fulfill constraints, such as IFEval (Zhou et al., 2023), INFOBench (Qin et al., 2024), and FollowBench (Jiang et al., 2023). These datasets focus on evaluating models’ abilities to handle compositional constraints, and have demonstrated promising advancements.
@@
Our work hypothesizes that requirements can be central to human-LLM interactions, as we have observed the LLM proficiency in following instructions. However, existing requirement datasets are often synthetic, containing prompts that are intentionally tricky or unrealistic, with requirements generated automatically via templates and limited to certain categories (e.g., sentence or paragraph length constraints, format constraints like output in JSON). While useful for identifying LLM-hard problems for requirement-focused training, these datasets do not fully capture how LLMs respond to real-world user prompts. Human prompts may have more diverse requirements across various tasks and less standardized language — a gap we seek to fulfill in this paper.
@@
Building on this foundation of instruction following, prompt optimizers have been proposed to improve the performance of frozen LLMs on user-defined downstream tasks by automatically refining user-written prompts. Automation ranges from simple adjustments like adding role-playing or chain-of-thought prompting (Wei et al., 2022) (e.g., Prompt Maker in Figure 2) to more advanced methods like TextGrad (Yuksekgonul et al., 2024), which searches for optimal wording, and DSPy (Singhvi et al., 2023), which extends the search space to few-shot prompts and various prompting techniques. While these sophisticated methods make more tailored changes, they are harder to apply and often require access to labeled datasets. Despite being in early stages, these optimizers aim to relieve humans from exhausting all possibilities in the natural language prompting space, as automated experiments with semantically-preserving edits can be more effective. We share their vision and foresee a future where optimizers play a crucial role in human-LLM interaction.
2.3. Requirements in Programming and Software Engineering

### v2

# What Should We Engineer in Prompts? Training Humans in Requirement-Driven LLM Use

@@ Abstract.
Prompting LLMs for complex tasks (e.g., building a trip advisor chatbot) needs humans to clearly articulate customized requirements (e.g., “start the response with a tl;dr”). However, existing prompt engineering instructions often lack focused training on requirement articulation and instead tend to emphasize increasingly automatable strategies (e.g., tricks like adding role-plays and “think step-by-step”). To address the gap, we introduce Requirement-Oriented Prompt Engineering (ROPE), a paradigm that focuses human attention on generating clear, complete requirements during prompting. We implement ROPE through an assessment and training suite that provides deliberate practice with LLM-generated feedback. In a randomized controlled experiment with 30 novices, ROPE significantly outperforms conventional prompt engineering training (20% vs. 1% gains), a gap that automatic prompt optimization cannot close. Furthermore, we demonstrate a direct correlation between the quality of input requirements and LLM outputs. Our work paves the way to empower more end-users to build complex LLM applications.
LLM, human-AI interaction, prompt engineering, requirement engineering, end-user programming
@@
Figure 1. We propose Requirement-Oriented Prompt Engineering (ROPE) and contribute a training and assessment suite for ROPE. (A) ROPE training helps learners write effective prompt programs by providing automated feedback on adding and clarifying requirements. (B) In our pre-post assessments, we ask learners to write prompts to develop customized applications (e.g., Trip Advisor via a prompt in C) and assess learners’ prompt quality on both requirement quality and LLM output quality. (D) In a pre-post randomized experiment, ROPE training significantly improves learners’ prompt quality, compared to conventional prompt engineering training.
@@ 1. Introduction
General-purpose AIs like large language models (LLMs) have evolved from simple next-word predictors (Brown, 2020) to powerful assistants capable of fulfilling complex user needs (Qin et al., 2024; Ouyang et al., 2022). This improvement in LLM’s ability to follow instructions has encouraged users to delegate increasingly intricate tasks to these models. In the early days of prompt engineering, users primarily focused on refining the wording of simple instructions to improve LLM output quality (Jiang et al., 2020; Shin et al., 2020). Today, prompts resemble detailed “essays” that define LLM behaviors. Rather than one-off small requests, these prompts start to power the generation of complex applications, and everyday users can write complex prompts to either generate code for softwares (Wodecki, 2024; Yang et al., 2024a) or tailor general-purpose LLMs into special purpose LLM applications (Zhang et al., 2024). For instance, an LLM application (or a GPTs) like Trip Advisor  [^1] https://chatgpt.com/g/g-K38J9feSL-trip-advisor can be built solely using prompts similar to Figure 1C. These LLM applications function similarly to traditional software and are accessible through LLM App Stores like GPT Store [^2] https://openai.com/index/introducing-the-gpt-store/ and FlowGPT [^3] https://flowgpt.com/, where they can be reused with a single click. In just a few months, over 30,000 “LLM app developers” have created millions of GPT apps, with the most popular ones being used more than 5 million times (Zhang et al., 2024). This trend signals a future of end-user programming through natural language, where anyone can create reusable programs just by writing prompts (Feng et al., 2024).
@@
The future of writing prompt programs looks promising, but status quo prompt engineering remains a challenge. Various studies show that optimizing LLM outputs requires well-rounded prompts, involving fluent writing, clear instructions, personas, formats, and effective adaption of prompting techniques like Chain-of-Thought (Bsharat et al., 2023; Meskó, 2023; Marvin et al., 2024; Schmidt et al., 2024). These complexities often confuse novices, causing them to make ad hoc revisions to prompts without a clear understanding of what needs improvement (Zamfirescu-Pereira et al., 2023). Recognizing the challenge, the NLP community is developing prompt optimizers to automate some prompt refinement steps, e.g., incorporating role-plays, applying effective techniques, structuring prompts, and enhancing text fluency (e.g., Figure 2, discussions in Section 2.2).
@@
The rise of optimizers makes us think: What aspects of prompt engineering should humans still master? We argue that the remaining part of manual prompt creation is akin to the core step before engineering in a design process: requirement specification. A requirement is a description of user goals; it defines and documents how a system should behave, including expected inputs and outputs (Lamsweerde, 2009) (e.g., ‘‘Start the response with a tl;dr’’ for Trip Advisor in Figure 1B). Requirements are the backbone of program functionalities, and user-specified requirements in prompts (e.g., Figure 2) are the driving forces behind complex LLM adaptations. Customized requirements are hard to predict, and thus difficult to fully automate and must be specified by humans. We argue that in a future where prompt programs are common, individuals should develop proficiency in requirement articulation. In other words, we need to shift our focus towards Requirement-Oriented Prompt Engineering, a novel paradigm that we call “ROPE.
@@
Articulating clear and complete requirements is a known challenge in fields such as product design and software engineering (Liu et al., 2023b; Nam et al., 2023; Wang and Zeng, 2009), and poor requirements often cause inefficiency, frustrations, and software failures (Zamfirescu-Pereira et al., 2023; Millett et al., 2007). This difficulty is amplified by the unpredictable nature of LLMs, which complicates controlling outputs (Arzberger et al., 2024; Chopra et al., 2023). Even experts need multiple iterations to refine requirements (Shankar et al., 2024a), and novices frequently make requirement mistakes (Babe et al., 2023). Recognizing both the importance and difficulty of this task, we pose the research question: Can we help end-users better instruct LLMs to achieve their goals through ROPE training?
@@
Figure 2. Automatically refine a user’s prompt for Trip Advisor using the optimizer Prompt Maker. Customized requirements still need to come from the human prompt, while the optimizer improves on non-requirement prompting aspects such as fluency, role plays, structures, etc.
@@
To answer this question, we develop the ROPE paradigm (Section 3), along with the requirement-focused training and assessments to (1) help users practice writing complete and correct requirements, and (2) deepen our understanding of how requirements affect the quality of LLM outputs. Specifically, we formalize tasks (Section 4.1) with complex requirements (e.g., game development), and design corresponding assessments (Section 4.2) that can measure the requirement quality, the LLM output quality, and their correlations. We also design and implement an interactive training mechanism (Section 4.3) that disentangles requirement articulation from other adjacent prompting aspects. Our system supports deliberate practice on requirement refinement by generating immediate feedback grounded on reference requirements.
@@
We conduct a randomized controlled user study (Section 5) with 30 prompting novices, comparing the requirements users write before and after training, and against a control group that received standard prompt engineering training. We confirm that our requirement-focused training is effective (Section 6): novices receiving this training significantly improved their requirement performances by two-fold in the pre-to-post test, achieved significantly more gains than conventional prompt engineering training, and learned to iterate on prompts in a more structured way towards requirements. Moreover, we confirm that requirements are indeed important for the LLM generation: There is a strong positive correlation between the requirement qualities and the LLM output qualities. Importantly, the training gains are complementary to potential improvements from optimizers: Using a popular prompt optimizer, we observe that optimizers can indeed improve requirement quality, but they cannot close the gap between the baseline and the training groups. We conclude with an in-depth discussion (Section 7) on future directions for human training and optimizer design to enhance human-LLM complementarity in ROPE.
In summary, we make the following contributions:
@@
As natural language prompt programs become more widespread, everyone needs the skill to write good requirements. We believe that with our ROPE paradigm, LLMs can empower anyone to program reusable AI assistants. We open-source our training system, supplemental study materials, and user prompts at https://github.com/mqo00/rope/.
@@ 2. Related Works
We review three key areas relevant to our work: prompt engineering, prompt optimization in natural language processing, and requirements engineering in end-user software engineering. We highlight a significant gap in the literature: the lack of requirement-oriented prompt engineering training for end users, which we propose to address in this paper.
@@ 2.1. Current Prompt Engineering Practices and Challenges
Prompt engineering (PE) has been essential for crafting effective input instructions to guide LLMs toward generating desired outputs (Liu et al., 2021; Bach et al., 2022; Zamfirescu-Pereira et al., 2023). However, the non-deterministic nature of LLMs makes it challenging for humans to predict LLMs’ behavior across prompts (Ouyang et al., 2024; Denny et al., 2024; Sun et al., 2022). This leads to wasted time on unproductive strategies, such as trivial wording changes (Feldman and Anderson, 2024; Denny et al., 2024; Zamfirescu-Pereira et al., 2023). While some challenges can be mitigated through automation (see Section 2.2), e.g., automating word choices, human requirements (some also refer to as constitutions or principles in literature (Petridis et al., 2024)) remain crucial for customized or specialized tasks (Liu et al., 2021). We define these tasks as LLM-hard, as a simple prompt cannot produce a satisfactory response. Developing reusable prompts for customized chatbots or GPTs (Bach et al., 2022; Zamfirescu-Pereira et al., 2023) is a common LLM-hard task. While a prompt is an input to LLMs, a reusable prompt is a set of instructions for recurring goals, essentially functioning as a prompt program. As more users create reusable prompts, the ability to write good prompt programs with clear requirements becomes increasingly important.
@@
However, prompt creation can be complex and inaccessible to non-experts who lack technical knowledge of LLMs, highlighting the need for more end-user prompt engineering (EUPE) scaffolding and training (Zamfirescu-Pereira et al., 2023). Various strategies have been developed to support PE, such as prompt technique catalog (Bsharat et al., 2023; Liu and Chilton, 2022; Wang et al., 2024a), chat-based tools like GPT-builder, [^4] https://chatgpt.com/create and toolkit for orchestrating complex tasks (Chase, 2022; Khattab et al., 2023). Each tool has trade-offs, balancing flexibility, precision, and technicality. For example, chat-based tools may ask narrow clarifications without fully understanding the context, leaving users to self-refine requirements when LLM responses are open-ended (Lahiri et al., 2022; Qian et al., 2024).
@@
Existing PE training emphasizes the importance of explicitly stating requirements within prompts (Meskó, 2023; Schmidt et al., 2024; Santana, 2024), yet there is still a lack of focused training on requirement articulation skills for everyday users. Without training, non-experts often make mistakes such as missing requirements and writing conflicting requirements in prompts (Feldman and Anderson, 2024; Nam et al., 2023; Prather et al., 2024; Mahdavi Goloujeh et al., 2024). Even experts need many iterations to improve their requirements for complicated LLM tasks (Shankar et al., 2024a). However, training novices on PE is difficult; for instance, novice programmers may not improve at prompting within a 75 minute study when provided with test cases and generated code as feedback (Nguyen et al., 2024). We aim to address this gap by proposing a ROPE training for end-users to improve their requirements engineering skills in the context of prompt creation.
@@ 2.2. Instruction Following for Foundation Models
Optimizing LLM performance on customized tasks typically follows two approaches: improving the models directly, or refining prompts with models fixed. The former has enabled the versatility of models like GPT-4, [^5] https://openai.com/index/gpt-4/ Gemini, [^6] https://gemini.google.com/ and LLaMA 3, [^7] https://www.llama.com/ which are fine-tuned through instruction tuning (Ouyang et al., 2022) and further aligned with human preferences using Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017). For the latter, the NLP community has been exploring LLMs’ capabilities to follow requirements (or “constraints”). Various datasets have been proposed to assess whether LLMs can fulfill constraints, such as IFEval (Zhou et al., 2023), INFOBench (Qin et al., 2024), and FollowBench (Jiang et al., 2024). These datasets focus on evaluating models’ abilities to handle compositional constraints, and have demonstrated promising advancements.
@@
Our work hypothesizes that requirements can be central to human-LLM interactions, as we have observed the LLM proficiency in following instructions. However, existing requirement datasets are often synthetic, containing prompts that are intentionally tricky or unrealistic, with requirements generated automatically via templates and limited to certain categories (e.g., sentence or paragraph length constraints, format constraints like output in JSON). While useful for identifying LLM-hard problems for requirement-focused training, these datasets do not fully capture how LLMs respond to real-world user prompts. Human prompts may have more diverse requirements across various tasks and less standardized language — a gap we seek to fill in this paper.
@@
Building on this foundation of instruction following, prompt optimizers have been proposed to improve the performance of frozen LLMs on user-defined downstream tasks by automatically refining user-written prompts. Automation ranges from simple adjustments like adding role-playing or chain-of-thought prompting (Wei et al., 2022) (e.g., Prompt Maker in Figure 2) to more advanced methods like TextGrad (Yuksekgonul et al., 2024), which searches for optimal wording, PromptCharm (Wang et al., 2024a), which enables users to adjust diffusion models’ attention to keywords in prompts, and DSPy (Khattab et al., 2023), which extends the search space to few-shot prompts and various prompting techniques. While these sophisticated methods make more tailored changes, they are harder to apply and often require access to labeled datasets. Despite being in early stages, these optimizers aim to relieve humans from exhausting all possibilities in the natural language prompting space, as automated experiments with semantically-preserving edits can be more effective. We share their vision and foresee a future where optimizers play a crucial role in human-LLM interaction.
2.3. Requirements in Software Engineering and Design